# Low Regularity of Self-Similar Solutions of Two-Dimensional Riemann problems with Shocks for the Isentropic Euler system

### v1

@@ 4.1. Lower Regularity of the Regular Shock Reflection Solutions for the Isentropic Euler System
Now we give the definition of a global regular shock reflection solution. The intuition for the definition is the following: The regular shock reflection solution is an entropy solution of the regular shock reflection problem in the sense of Definition 4.2 which has the structure shown in Figs. 4.24.2, where $(\rho,{\mathbf{v}})$ coincides with those of states $(0)$, $(1)$, and $(2)$ in their respective regions. As we discussed above, the necessary condition is the existence of state $(2)$, which means that the wedge angle satisfies $\theta_{\rm w}\in(\theta_{\rm w}^{\rm d},\frac{\pi}{2})$. Moreover, it is expected that the solution is relatively regular in $\Omega$. However, we show below that it is not possible that $(\rho,{\mathbf{v}})\in H^{1}(\Omega)$. On the other hand, from the physical/computational experiments and the theoretical results in the case of potential flow, it is expected that $\Gamma_{\rm shock}$ is a smooth curve and $(\rho,{\mathbf{v}})$ is smooth near and up to $\Gamma_{\rm shock}\cup\Gamma_{\rm sonic}$ in $\Omega$. Moreover, the regularity discussed is expected for the shock reflection-diffraction configuration extended to $\{\xi_{1}<0\}$ by the even reflection about the $\xi_{1}$–axis (since this is the original shock reflection-diffraction configuration). In particular, the extended shock curve $\Gamma_{\text{\rm shock}}^{\rm ext}$ is smooth, which shows that $\Gamma_{\rm shock}$ must be orthogonal to $\Gamma_{\rm sym}$ at $P_{2}$. Then, noting that the pseudo-velocity of state $(1)$ is ${\mathbf{v}}_{1}({\boldsymbol{\xi}})=(u_{1}^{(1)},0)-{\boldsymbol{\xi}}$ with $u_{1}^{(1)}>0$ and $\{P_{2}\}=\overline{\Gamma}_{\rm shock}\cap\overline{\Gamma}_{\rm sym}\subset\{\xi_{1}<0\}$, it follows that vector ${\mathbf{v}}_{1}$ on $\Gamma_{\rm shock}$ near $P_{2}$ points into $\Omega$. It is then expected that this holds on the whole shock $\Gamma_{\rm shock}$, unless the jumps of the velocity and the density across the reflected-diffracted shock are degenerate at some points (which are not expected). Then it follows from (2.10) for $\Gamma_{\rm shock}$ that ${\mathbf{v}}$ on $\Gamma_{\rm shock}$ from the $\Omega$–side also points into $\Omega$. On the sonic arc (for the supersonic reflection), the jump of the velocity is not expected, and velocity ${\mathbf{v}}_{2}$ of state $(2)$ on $\Gamma_{\rm sonic}$ points into $\Omega$ (as ${\mathbf{v}}_{2}$ points along the radial direction of the sonic circle of state $(2)$ towards its center $O_{2}\in P_{3}P_{4}$). Therefore, it is expected that
@@
If state $(2)$ for $\theta_{\rm w}$ is supersonic at $P_{0}$, i.e., $|{\mathbf{v}}_{2}(P_{0})|>c_{2}$, the solution has the supersonic reflection structure as on Fig. 4.2. If state $(2)$ for $\theta_{\rm w}$ is subsonic or sonic at $P_{0}$, i.e., $|{\mathbf{v}}_{2}(P_{0})|\leq c_{2}$, the solution has the subsonic reflection structure as on Fig. 4.2. More specifically,
@@
    $(\rho,{\mathbf{v}})$ is continuous in $\overline{\Omega}\cap{\mathcal{N}}_{r}(\Gamma_{\rm shock}\cup\Gamma_{\rm sonic})$ for the supersonic or sonic reflections, and in $\overline{\Omega}\cap{\mathcal{N}}_{r}(\Gamma_{\rm shock})$ for the subsonic reflection, for some $r>0$.
@@
    $(\rho,{\mathbf{v}})=\begin{cases}(\rho_{0},{\mathbf{v}}_{0})\qquad\mbox{for $\xi_{1}>\xi_{1}^{0}$ and $\xi_{2}>\xi_{1}\tan\theta_{\rm w}$},\\[2.84526pt] (\rho_{1},{\mathbf{v}}_{1})\qquad\mbox{for $\xi_{1}<\xi_{1}^{0}$ and above curve $P_{0}P_{1}P_{2}$},\\[2.84526pt] (\rho_{2},{\mathbf{v}}_{2})\qquad\mbox{in $P_{0}P_{1}P_{4}$},\end{cases}$
@@
    $(\rho,{\mathbf{v}})=\begin{cases}(\rho_{0},{\mathbf{v}}_{0})\qquad\mbox{for $\xi_{1}>\xi_{1}^{0}$ and $\xi_{2}>\xi_{1}\tan\theta_{\rm w}$},\\[2.84526pt] (\rho_{1},{\mathbf{v}}_{1})\qquad\mbox{for $\xi_{1}<\xi_{1}^{0}$ and above curve $P_{0}P_{2}$},\end{cases}$
@@
    The flow is pseudo-subsonic in $\Omega$ on and near $\Gamma_{\rm shock}$, except the except the sonic point $P_{1}$ for the supersonic or sonic reflections, i.e., $|{\mathbf{v}}|<c$ on $\overline{\Gamma}_{\rm shock}\setminus\{P_{1}\}$ for supersonic and sonic reflections, and on $\overline{\Gamma}_{\rm shock}$ for strictly subsonic reflections, where ${\mathbf{v}}$ on $\Gamma_{\rm shock}$ is computed from the $\Omega$–side.
@@
For the supersonic reflections, it follows from conditions (ii)–(iii) of Definition 4.3 and the Rankine-Hugoniot conditions that $(\rho,{\mathbf{v}})$ are continuous across the sonic arc $P_{1}P_{4}=\overline{\Gamma_{\rm sonic}}$:
@@
Let $(\rho,{\mathbf{v}})$ be a regular shock reflection solution in the sense of Definition 4.3. Assume that $(\rho,{\mathbf{v}})$ satisfy the following:
@@
Then it is not possible that $(\rho,{\mathbf{v}})\in H^{1}(\Omega)$.
@@
Next, for the non-symmetric regular shock reflection problem, based on the argument near the reflection point $P_{0}$ for the symmetric case above, there are four configurations depending on whether the solution at $P_{0}$ and $P_{1}$ is subsonic or supersonic. As two examples of them, Fig. 4.4 and Fig. 4.4 exhibit the structures of two supersonic regular reflections and two subsonic regular reflections, respectively. In light of Definition 4.3, we define the regular reflection solution for the non-symmetric case. Let the wedge angle $\theta_{\mathrm{w}}=\theta_{\mathrm{w}}^{1}+\theta_{\mathrm{w}}^{2}$ with $\theta_{\mathrm{w}}^{i}\in(\theta_{\mathrm{w}}^{\mathrm{d}},\frac{\pi}{2})$, where $\theta_{\mathrm{w}}^{i}$ is the angle between $\Gamma_{\mathrm{w}}^{i}$ and the $\xi_{1}$–axis for $i=1,2$. Let
@@
Figure 4.3. Non-symmetric supersonic regular reflection
Figure 4.4. Non-symmetric subsonic regular reflection
@@
$(\bar{\rho},\bar{{\mathbf{v}}})=\begin{cases}(\rho_{0},{\mathbf{v}}_{0})\qquad\mbox{for $\xi_{1}>\xi_{1}^{0}$},\\[2.84526pt] (\rho_{1},{\mathbf{v}}_{1})\qquad\mbox{for $\xi_{1}<\xi_{1}^{0}$}.\end{cases}$
@@
    If state $(2)$ at $P_{0}$ and state $(3)$ at $P_{1}$ are both supersonic, i.e., $|{\mathbf{v}}_{2}(P_{0})|>c_{2}$ and $|{\mathbf{v}}_{2}(P_{1})|>c_{3}$, the solution has the supersonic reflection structure as in Fig. 4.4 at both points $P_{0}$ and $P_{1}$. If state $(2)$ at $P_{0}$ and state $(3)$ at $P_{1}$ are both subsonic or sonic, i.e., $|{\mathbf{v}}_{2}(P_{0})|\leq c_{2}$ and $|{\mathbf{v}}_{2}(P_{1})|\leq c_{3}$, the solution has the subsonic reflection structure as in Fig. 4.4.
@@
    $(\rho,{\mathbf{v}})=\begin{cases}(\rho_{0},{\mathbf{v}}_{0})\qquad\mbox{for}\,\,\,\xi_{1}>\xi_{1}^{0},\,\,\,\xi_{2}>\xi_{1}\tan\theta_{\rm w}^{1},\\[2.84526pt] (\rho_{0},{\mathbf{v}}_{0})\qquad\mbox{for}\,\,\,\xi_{1}>\xi_{1}^{0},\,\,\,\xi_{2}<-\xi_{1}\tan\theta_{\rm w}^{2},\\[2.84526pt] (\rho_{1},{\mathbf{v}}_{1})\qquad\mbox{for}\,\,\,\xi_{1}<\xi_{1}^{0},\,\,\,\mbox{and the left to curve }\;P_{0}P_{1},\\[2.84526pt] (\rho_{2},{\mathbf{v}}_{2})\qquad\mbox{in}\,\,\,\Lambda_{2}\,\,\,\mbox{if state $(2)$ at $P_{0}$ is supersonic},\\[2.84526pt] (\rho_{3},{\mathbf{v}}_{3})\qquad\mbox{in}\,\,\,\Lambda_{3}\,\,\,\mbox{if state $(3)$ at $P_{1}$ is supersonic},\end{cases}$
    where $\xi_{1}^{0}>0$ is the location of the incident shock $S_{0}$ on the self-similar plane, and
    $\lim_{{\boldsymbol{\xi}}\in\Omega,\;{\boldsymbol{\xi}}\to P_{0}}(\rho,{\mathbf{v}})({\boldsymbol{\xi}})=(\rho_{2},{\mathbf{v}}_{2})(P_{0})\qquad\mbox{if state $(2)$ at $P_{0}$ is subsonic or sonic},$
    $\lim_{{\boldsymbol{\xi}}\in\Omega,\;{\boldsymbol{\xi}}\to P_{1}}(\rho,{\mathbf{v}})({\boldsymbol{\xi}})=(\rho_{3},{\mathbf{v}}_{3})(P_{1})\qquad\mbox{if state $(3)$ at $P_{1}$ is subsonic or sonic}.$
@@
For the supersonic reflections, it follows from conditions (ii)–(iv) of Definition 4.7 and the Rankine-Hugoniot conditions that $(\rho,{\mathbf{v}})$ are continuous across the sonic arcs $\Gamma_{\rm sonic}^{2}\cup\Gamma_{\rm sonic}^{3}$:
@@
Let $(\rho,{\mathbf{v}})$ be a non-symmetric regular shock reflection solution in the sense of Definition 4.7. Assume that $(\rho,{\mathbf{v}})$ satisfy the following:
@@
Then it is not possible that $(\rho,{\mathbf{v}})\in H^{1}(\Omega)$.
@@
It suffices to check that the assumptions of Theorem 3.2 are satisfied. We divide the proof into four steps.
1. Since the cases with one supersonic reflection and one subsonic reflection can be treated similarly, we focus only on the non-symmetric supersonic reflection solution (see Fig. 4.4) and the non-symmetric subsonic reflection solution (see Fig. 4.4) in the proof below.
@@
${\mathbf{v}}\cdot{\boldsymbol{\nu}}=-c_{2}<0\qquad\mbox{on $\Gamma_{\rm sonic}^{2}$ and ${\mathbf{v}}\cdot{\boldsymbol{\nu}}=-c_{3}<0$ on $\Gamma_{\rm sonic}^{3}$}.$
@@
Condition (a) is satisfied since $\Lambda_{1}$ is the region of state $(1)$ as shown in Figs. 4.44.4 above, and $\Gamma_{\rm shock}=\partial\Lambda_{1}\cap\partial\Omega$ in both the supersonic and subsonic cases.
@@
Consider first the case of reflections that are supersonic at $P_{0}$, i.e., $|{\mathbf{v}}_{2}(P_{0})|>c_{2}$; see Fig. 4.4. Then we need to show that $({\mathbf{v}}\cdot{\boldsymbol{\tau}}_{\rm shock})(P_{2})\neq 0$. Assume this is not true, then $({\mathbf{v}}\cdot{\boldsymbol{\tau}}_{\rm shock})(P_{2})=0$. Note that ${\boldsymbol{\tau}}_{\rm shock}(P_{2})={\boldsymbol{\tau}}_{S_{1}}$ and ${\mathbf{v}}_{|\Omega}(P_{2})={\mathbf{v}}_{2}(P_{2})$ by Definition 4.7(ii, iv) so that ${\mathbf{v}}_{2}(P_{2})\cdot{\boldsymbol{\tau}}_{S_{1}}=0$ and
@@
In the case of reflections that are sonic at $P_{0}$, i.e., $|{\mathbf{v}}_{2}(P_{0})|=c_{2}$ (see Fig. 4.4), the argument is the same as above with only notational change: we use point $P_{0}$ here instead of point $P_{2}$.
It remains to consider the case of reflections that are subsonic at $P_{0}$, i.e., $|{\mathbf{v}}_{2}(P_{0})|<c_{2}$; see Fig. 4.4. Assume that $({\mathbf{v}}\cdot\boldsymbol{\tau}_{\rm shock})(P_{0})=0$. Recall that $(\rho,{\mathbf{v}})(P_{0})=(\rho_{2},{\mathbf{v}}_{2})(P_{0})$ for the subsonic reflection and $S_{1}$ is tangent to $\Gamma_{\rm shock}$ at $P_{0}$, as we have shown above. Thus, from $({\mathbf{v}}\cdot\boldsymbol{\tau}_{\rm shock})(P_{0})=0$, we obtain that $({\mathbf{v}}_{2}\cdot\boldsymbol{\tau}_{S_{1})}(P_{0})=0$. Since $S_{1}$ is the line shock between states $(1)$ and $(2)$, the last equality implies by (2.10) that $({\mathbf{v}}_{1}\cdot\boldsymbol{\tau}_{S_{1})}(P_{0})=0$, so that line $L$ through centers $O_{1}=(u_{1}^{(1)},0)$ and $O_{2}=(u_{1}^{(2)},u_{2}^{(2)})$ of states $(1)$ and $(2)$ (which is orthogonal to $S_{1}$) intersects $S_{1}$ at $P_{0}$. Since $P_{0}\in\Gamma_{\rm wedge}$, $O_{2}\in L\cap\Gamma_{\rm wedge}$ (where $O_{2}\in\Gamma_{\rm wedge}$ because ${\mathbf{v}}_{2}\cdot{\boldsymbol{\nu}}=0$ on $\Gamma_{\rm wedge}$), and $L\cap\Gamma_{\rm wedge}=\{P_{0}\}$, it follows that $O_{2}=P_{0}$ so that ${\mathbf{v}}_{2}(P_{0})=\bf{0}$. Since ${\mathbf{v}}_{2}(P_{0})=\bf{0}$, it follows from the Rankine-Hugoniot conditions (2.9)–(2.10) on $S_{1}$ between states $(1)$ and $(2)$ that ${\mathbf{v}}_{1}(P_{0})=\bf{0}$, that is, $O_{1}=P_{0}$. However, this is not true since $O_{1}=(u_{1}^{(1)},0)$ for $u_{1}^{(1)}>0$, while $P_{0}=(l\cos\theta_{\rm w}^{1},l\sin\theta_{\rm w}^{1})$ for some $l>0$. This contradiction shows that $({\mathbf{v}}\cdot\boldsymbol{\tau}_{\rm shock})(P_{0})\neq 0$, i.e., condition (c) of Lemma 4.1 holds for the subsonic reflection at $P_{0}$.

### v2

@@ 4.1. Lower Regularity of the Regular Shock Reflection Solutions for the Isentropic Euler System
Now we give the definition of a global regular shock reflection solution. The intuition for the definition is the following: The regular shock reflection solution is an entropy solution of the regular shock reflection problem in the sense of Definition 4.2 which has the structure shown in Figs. 33, where $(\rho,{\mathbf{v}})$ coincides with those of states $(0)$, $(1)$, and $(2)$ in their respective regions. As we discussed above, the necessary condition is the existence of state $(2)$, which means that the wedge angle satisfies $\theta_{\rm w}\in(\theta_{\rm w}^{\rm d},\frac{\pi}{2})$. Moreover, it is expected that the solution is relatively regular in $\Omega$. However, we show below that it is not possible that ${\mathbf{v}}\in H^{1}(\Omega)$. On the other hand, from the physical/computational experiments and the theoretical results in the case of potential flow, it is expected that $\Gamma_{\rm shock}$ is a smooth curve and $\mathbf{v}$} is smooth near and up to $\Gamma_{\rm shock}\cup\Gamma_{\rm sonic}$ in $\Omega$. Moreover, the regularity discussed is expected for the shock reflection-diffraction configuration extended to $\{\xi_{1}<0\}$ by the even reflection about the $\xi_{1}$–axis (since this is the original shock reflection-diffraction configuration). In particular, the extended shock curve $\Gamma_{\text{\rm shock}}^{\rm ext}$ is smooth, which shows that $\Gamma_{\rm shock}$ must be orthogonal to $\Gamma_{\rm sym}$ at $P_{2}$. Then, noting that the pseudo-velocity of state $(1)$ is ${\mathbf{v}}_{1}({\boldsymbol{\xi}})=(u_{1}^{(1)},0)-{\boldsymbol{\xi}}$ with $u_{1}^{(1)}>0$ and $\{P_{2}\}=\overline{\Gamma_{\rm shock}}\cap\overline{\Gamma_{\rm sym}}\subset\{\xi_{1}<0\}$, it follows that vector ${\mathbf{v}}_{1}$ on $\Gamma_{\rm shock}$ near $P_{2}$ points into $\Omega$. It is then expected that this holds on the whole shock $\Gamma_{\rm shock}$, unless the jumps of the velocity and the density across the reflected-diffracted shock are degenerate at some points (which are not expected). Then it follows from (2.10) for $\Gamma_{\rm shock}$ that $\mathbf{v}$} on $\Gamma_{\rm shock}$ from the $\Omega$–side also points into $\Omega$. On the sonic arc (for the supersonic reflection), the jump of the velocity is not expected, and velocity ${\mathbf{v}}_{2}$ of state $(2)$ on $\Gamma_{\rm sonic}$ points into $\Omega$ (as ${\mathbf{v}}_{2}$ points along the radial direction of the sonic circle of state $(2)$ towards its center $O_{2}\in P_{3}P_{4}$). Therefore, it is expected that
@@
If state $(2)$ for $\theta_{\rm w}$ is supersonic at $P_{0}$, i.e., $|{\mathbf{v}}_{2}(P_{0})|>c_{2}$, the solution has the supersonic reflection structure as on Fig. 3. If state $(2)$ for $\theta_{\rm w}$ is subsonic or sonic at $P_{0}$, i.e., $|{\mathbf{v}}_{2}(P_{0})|\leq c_{2}$, the solution has the subsonic reflection structure as on Fig. 3. More specifically,
@@
    $(\rho,{\mathbf{v}})$ is continuous in $\overline{\Omega}\cap{\mathcal{N}}_{r}(\Gamma_{\rm shock}\cup\Gamma_{\rm sonic})$ for the supersonic reflection, and in $\overline{\Omega}\cap{\mathcal{N}}_{r}(\Gamma_{\rm shock})$ for the subsonic or sonic reflection, for some $r>0$.
@@
[…]
@@
[…]
@@
    The flow is pseudo-subsonic in $\Omega$ on and near $\Gamma_{\rm shock}$, except for the sonic point $P_{1}$ for the supersonic or sonic reflections, i.e., $|{\mathbf{v}}|<c$ on $\overline{\Gamma_{\rm shock}}\setminus\{P_{1}\}$ for supersonic and sonic reflections, and on $\overline{\Gamma_{\rm shock}}$ for strictly subsonic reflections, where $\mathbf{v}$} on $\Gamma_{\rm shock}$ is computed from the $\Omega$–side.
@@
For the supersonic reflections, it follows from conditions (ii)–(iii) of Definition 4.3 and the Rankine-Hugoniot conditions that $(\rho,{\mathbf{v}})$ is continuous across the sonic arc $P_{1}P_{4}=\overline{\Gamma_{\rm sonic}}$:
@@
Let $(\rho,{\mathbf{v}})$ be a regular shock reflection solution in the sense of Definition 4.3. Assume that $(\rho,{\mathbf{v}})$ satisfies the following:
@@
Then ${\mathbf{v}}\notin H^{1}(\Omega)$.
@@
Next, for the non-symmetric regular shock reflection problem, based on the argument near the reflection point $P_{0}$ for the symmetric case above, there are four configurations depending on whether the solution at $P_{0}$ and $P_{1}$ is subsonic or supersonic. As two examples of them, Fig. 5 and Fig. 5 exhibit the structures of two supersonic regular reflections and two subsonic regular reflections, respectively. In light of Definition 4.3, we define the regular reflection solution for the non-symmetric case. Let the wedge angle $\theta_{\mathrm{w}}=\theta_{\mathrm{w}}^{1}+\theta_{\mathrm{w}}^{2}$ with $\theta_{\mathrm{w}}^{i}\in(\theta_{\mathrm{w}}^{\mathrm{d}},\frac{\pi}{2})$, where $\theta_{\mathrm{w}}^{i}$ is the angle between $\Gamma_{\mathrm{w}}^{i}$ and the $\xi_{1}$–axis for $i=1,2$. Let
@@
Figure 4. Non-symmetric supersonic regular reflection
Figure 5. Non-symmetric subsonic regular reflection
@@
[…]
@@
    If state $(2)$ at $P_{0}$ and state $(3)$ at $P_{1}$ are both supersonic, i.e., $|{\mathbf{v}}_{2}(P_{0})|>c_{2}$ and $|{\mathbf{v}}_{2}(P_{1})|>c_{3}$, the solution has the supersonic reflection structure as in Fig. 5 at both points $P_{0}$ and $P_{1}$. If state $(2)$ at $P_{0}$ and state $(3)$ at $P_{1}$ are both subsonic or sonic, i.e., $|{\mathbf{v}}_{2}(P_{0})|\leq c_{2}$ and $|{\mathbf{v}}_{2}(P_{1})|\leq c_{3}$, the solution has the subsonic reflection structure as in Fig. 5.
@@
[…]
    where $\xi_{1}^{0}>0$ is the location of the incident shock $S_{0}$ on the self-similar plane, and limξΩ, ξP0(ρ, v)(ξ)=(ρ2, v2)(P0)  if state $(2)$ at $P_{0}$ is subsonic or sonic, limξΩ, ξP1(ρ, v)(ξ)=(ρ3, v3)(P1)  if state $(3)$ at $P_{1}$ is subsonic or sonic.
@@
For the supersonic reflections, it follows from conditions (ii)–(iv) of Definition 4.7 and the Rankine-Hugoniot conditions that $(\rho,{\mathbf{v}})$ is continuous across the sonic arcs $\Gamma_{\rm sonic}^{2}\cup\Gamma_{\rm sonic}^{3}$:
@@
Let $(\rho,{\mathbf{v}})$ be a non-symmetric regular shock reflection solution in the sense of Definition 4.7. Assume that $(\rho,{\mathbf{v}})$ satisfies the following:
@@
Then ${\mathbf{v}}\notin H^{1}(\Omega)$.
@@
It suffices to check that the assumptions of Theorem 3.2 are satisfied. We divide the proof into four steps:
1. Since the cases with one supersonic reflection and one subsonic reflection can be treated similarly, we focus only on the non-symmetric supersonic reflection solution (see Fig. 5) and the non-symmetric subsonic reflection solution (see Fig. 5) in the proof below.
@@
[…]
@@
Condition (a) is satisfied since $\Lambda_{1}$ is the region of state $(1)$ as shown in Figs. 55 above, and $\Gamma_{\rm shock}=\partial\Lambda_{1}\cap\partial\Omega$ in both the supersonic and subsonic cases.
@@
Consider first the case of reflections that are supersonic at $P_{0}$, i.e., $|{\mathbf{v}}_{2}(P_{0})|>c_{2}$; see Fig. 5. Then we need to show that $({\mathbf{v}}\cdot{\boldsymbol{\tau}}_{\rm shock})(P_{2})\neq 0$. Assume this is not true, then $({\mathbf{v}}\cdot{\boldsymbol{\tau}}_{\rm shock})(P_{2})=0$. Note that ${\boldsymbol{\tau}}_{\rm shock}(P_{2})={\boldsymbol{\tau}}_{S_{1}}$ and ${\mathbf{v}}_{|\Omega}(P_{2})={\mathbf{v}}_{2}(P_{2})$ by Definition 4.7(ii, iv) so that ${\mathbf{v}}_{2}(P_{2})\cdot{\boldsymbol{\tau}}_{S_{1}}=0$ and
@@
In the case of reflections that are sonic at $P_{0}$, i.e., $|{\mathbf{v}}_{2}(P_{0})|=c_{2}$ (see Fig. 5), the argument is the same as above with only notational change: we use point $P_{0}$ here instead of point $P_{2}$.
It remains to consider the case of reflections that are subsonic at $P_{0}$, i.e., $|{\mathbf{v}}_{2}(P_{0})|<c_{2}$; see Fig. 5. Assume that $({\mathbf{v}}\cdot\boldsymbol{\tau}_{\rm shock})(P_{0})=0$. Recall that $(\rho,{\mathbf{v}})(P_{0})=(\rho_{2},{\mathbf{v}}_{2})(P_{0})$ for the subsonic reflection and $S_{1}$ is tangent to $\Gamma_{\rm shock}$ at $P_{0}$, as we have shown above. Thus, from $({\mathbf{v}}\cdot\boldsymbol{\tau}_{\rm shock})(P_{0})=0$, we obtain that $({\mathbf{v}}_{2}\cdot\boldsymbol{\tau}_{S_{1})}(P_{0})=0$. Since $S_{1}$ is the line shock between states $(1)$ and $(2)$, the last equality implies by (2.10) that $({\mathbf{v}}_{1}\cdot\boldsymbol{\tau}_{S_{1})}(P_{0})=0$, so that line $L$ through centers $O_{1}=(u_{1}^{(1)},0)$ and $O_{2}=(u_{1}^{(2)},u_{2}^{(2)})$ of states $(1)$ and $(2)$ (which is orthogonal to $S_{1}$) intersects $S_{1}$ at $P_{0}$. Since $P_{0}\in\Gamma_{\rm wedge}$, $O_{2}\in L\cap\Gamma_{\rm wedge}$ (where $O_{2}\in\Gamma_{\rm wedge}$ because ${\mathbf{v}}_{2}\cdot{\boldsymbol{\nu}}=0$ on $\Gamma_{\rm wedge}$), and $L\cap\Gamma_{\rm wedge}=\{P_{0}\}$, it follows that $O_{2}=P_{0}$ so that ${\mathbf{v}}_{2}(P_{0})=\bf{0}$. Since ${\mathbf{v}}_{2}(P_{0})=\bf{0}$, it follows from the Rankine-Hugoniot conditions (2.9)–(2.10) on $S_{1}$ between states $(1)$ and $(2)$ that ${\mathbf{v}}_{1}(P_{0})=\bf{0}$, that is, $O_{1}=P_{0}$. However, this is not true since $O_{1}=(u_{1}^{(1)},0)$ for $u_{1}^{(1)}>0$, while $P_{0}=(l\cos\theta_{\rm w}^{1},l\sin\theta_{\rm w}^{1})$ for some $l>0$. This contradiction shows that $({\mathbf{v}}\cdot\boldsymbol{\tau}_{\rm shock})(P_{0})\neq 0$, i.e., condition (c) of Lemma 4.1 holds for the subsonic reflection at $P_{0}$.